Please help me solve this problem. A basketball player scored 100 points in a game. He took 60 shots. If each shot scored either one or two points, how many of each did he shoot? I need to see the work, because its really hard for me to understand. Thank you
Answer by richwmiller(17219)

a+b=60,
1*a+2*b=100
a=60-b
1*(60-b)+2*b=100
60-1b+2*b=100
1*b=40
b=40
a=60-b
a=20 at 1 point and b=40 at 2 points
check
20+40=60
1*a+2*b=100
1*20+2*40=100
20+80=100
100=100
ok
codetickv